# Optimal Deterministic Rendezvous in Labeled Lines

### v1

@@ 3.3 Computing Colored Ruling Sets on Paths, and Stopping Early
Our algorithm gradually computes a ruling set $S$ of the infinite line, in the sense that we provide an algorithm in which every node on the infinite line eventually decides whether it is part of the ruling set or not, and in any given round, the set of nodes that committed to their output and joined the set $S$ is a ruling set over the nodes that committed to their output. At a high level, Algorithm 3 partitions the nodes on the infinite line according to their IDs, and computes a ruling set over each set of nodes of this partition using Algorithm 2 (PathRulingSet) from the previous section. The computation done by the set of nodes $V_{i}=\{v\in V:\log^{*}\ell(v)=i\}$ is well defined as it is performed by a finite set of nodes, with bounded IDs. After $\Theta(R\cdot i)$ rounds, the nodes in $V_{i}$ have all finished their execution of PathRulingSet, as did previously nodes in the sets $V_{j}$, $j<i$. Once $V_{i}$ has computed a $V_{i}$-limited ruling set $S_{i}$, we use it to extend the current ruling set $S$, that only contains nodes from $V_{<i}$ this point. This is done greedily. We first have each node $v\in S_{i}$ join $S$ if and only if $S$ does not already contain a node at distance $<R$ from $v$ on the line. A second greedy extension slightly improves the covering distance. Finally, we assign a color to the new members of the ruling set.
@@
2:$S\leftarrow\emptyset$ current ruling set, $\chi:S\to[17]$ current coloring.
@@
4:     Compute a $V_{i}$-limited $(R,R-1)$-ruling set with PathRulingSet.
@@
7:     Let $S_{<i}$ be the value of $S$ at this stage. $S_{<i}$ is a $V_{<i}$-limited $(R,R-1)$-ruling set.
@@
10:     for $4$ iterations do
11:         For each $v\in V_{i}$, compute $b_{v}=\min_{u\in S}\operatorname{dist}(v,u)$. Let $U_{\geq R}=\{v\in U:b_{v}\geq R\}$.
12:         Add to $S$ all $v\in U_{\geq R}$ s.t. $\forall u\in N^{R-1}(v)\cap U_{\geq R},(b_{u}<b_{v}\vee(b_{u}=b_{v}\wedge\ell(v)>\ell(u)))$.
@@
Algorithm 3 achieves the following:
@@
    At termination, $v$ outputs whether it is part of the ruling set $S$ or not, and if part of $S$, it also outputs a color between $1$ and $17$. Nodes outside $S$ also know the members of $S$ at distance $<R$ together with their colors.
    For any round $T$, let $U_{T}$ be the set of nodes that have terminated up to that round, $S_{T}\subseteq U_{T}$ the nodes the set of nodes that joined $S$, and $\chi_{T}:S_{T}\to[17]$
@@
    $S_{T}$ is a $U_{T}$-limited $(R,R-1)$-ruling set, and $\chi_{T}$ is a proper coloring of $G^{9R-1}[S_{T}]$.
@@
We first show that Algorithm 3 (ES-ColPathRulingSet) constructs the claimed object, before bounding its complexity.
Properties of the Ruling Set
Recall that $S_{i}$, the output of PathRulingSet executed by nodes in $V_{i}$, is a $V_{i}$-limited $(R,R-1)$-ruling set. Consider the set $S_{<i}$, the content of $S$ before adding nodes from $V_{i}$ to it. Assume $S_{<i}$ to be a $V_{<i}$-limited $(R,R-1)$-ruling set. We show that the nodes from $V_{i}$ added to it make $S$ a $V_{\leq i}$-limited $(R,R-1)$-ruling set. By induction, starting from the trivial base case of $S_{<1}$, we get the claimed properties about our ruling set.
Every node $v\in V_{i}$ is at distance at most $R-1$ from a node $u\in S_{i}\subseteq V_{i}$. Consider the state of $S$ after adding $S^{\prime}_{i}=S_{i}\setminus N^{R-1}(S)$ to it. If $u$ was added to $S$, then $v$ remains at distance at most $R-1$ from a node in $S$. If $u$ was not added to $S$, there exists a node $w\in S_{<i}$ that joined $S$ earlier, such that $\operatorname{dist}(u,w)<R$. This implies that $v$ is at distance at most $\operatorname{dist}(v,u)+\operatorname{dist}(u,w)\leq 2R-2$ from a node in $S$. At this state, $S=S_{<i}\cup S^{\prime}_{i}$ is a $V_{\leq i}$-limited $(R,2R-2)$-ruling set. The $4$ iterations of the inner loop that add to $S$ some of the nodes in $V_{i}$ furthest away from the current set $S$ make $S$ a $V_{\leq i}$-limited $(R,R-1)$-ruling set, from similar arguments as used in the proof of Lemma 3.4.
@@
Complexity
Recall that PathRulingSet (Algorithm 2) with parameter distance $R\geq 1$ has a complexity of $O(R\log^{*}n)$ rounds when run by nodes of maximum ID $n$. Nodes in $V_{i}$ thus spend $O(R\cdot i)$ rounds computing $S_{i}$ at the beginning of the (parallel) loop in Algorithm 3.
For every $i\in\mathbb{N}^{*}$, let $T_{i}$ be the round in which nodes in $V_{i}$ terminate in Algorithm 3. Nodes in $V_{i}$ first compute $S_{i}$, before greedily adding nodes of $S_{i}$ to $S$. Before adding nodes of $S_{i}$ to $S$, all nodes in $V_{j},j<i$ must have terminated, which requires waiting until round $T_{i-1}$ (00 for the first set of nodes). After this amount of time, inserting the nodes of $S_{i}$ into $S$ only takes $R-1$ rounds, for each node in $S_{i}$ to learn which nodes in its $(R-1)$-neighborhood have joined $S$. The $4$ loop iterations in which additional nodes from $V_{i}$ are added to $S$ are also completed in $O(R)$. Finally, solving the list-coloring problem is done in $O(R\cdot i)$ rounds (Lemma 2.1), and each node learning the output of nodes within distance $R$ takes $O(R)$ rounds.
@@ 4.1 The Algorithm
In this section, we describe our deterministic rendezvous algorithm. We actually describe an algorithm working in the setting where agents achieve rendezvous when crossing the same edge in opposite directions. By Theorem 3.1, this implies an algorithm with similar asymptotic complexity for the setting where agents only detect each other if at the same node in a given round.
@@
The two basic building blocks of our algorithm are ZWalk and ES-ColPathRulingSet, respectively Algorithm 4 and Algorithm 3. ZWalk is simply a procedure in which the agent executing it scans the $12L+$ 1 nodes of the distance-$L$ neighborhood of the node it starts the ZWalk from, in $4L$ rounds. Note that while the agents do not have a common orientation, they have a sense of direction in that they know the port from which they enter a node. This allows the agents to keep walking in the same direction for successive steps, and to change direction, as desired. Most moves taken during our algorithm are done in executions of ZWalk. ES-ColPathRulingSet is the early-stopping algorithm for computing colored ruling sets presented in the previous section.
@@
Our algorithm (Algorithm 5) consists of a loop with a parameter $L$ that doubles between iterations. Inside the loop are two phases, which we call the discovery phase and the searching phase. In the discovery phase, an agent explore the distance-$L$ neighborhood around their starting position, roughly doubling how much of the network they know about. The agents use this information to compute a colored ruling set of parameter $R$, where $R\leq L/16$ is the largest power of $4$ s.t. the distance-$O(R)$ neighborhood of their starting location contains a ruling-set node. Of course, as the agents have different views, they do not necessarily end up computing a ruling-set with the same parameter $R$, and we denote by $R_{\alpha,L}$ and $R_{\beta,L}$ the distance of the ruling sets computed by agents $\alpha$ and $\beta$ in their phase $L$. The agents use their respective computed ruling set to search for one another, using procedure SearchingWalk (Algorithm 6), which constitutes the searching phase.
@@
6:     if $\exists R\in\{4^{j}\mid j\in\mathbb{N}\}$, $L\leq R/16$, s.t. $S_{R}\neq\emptyset$ then
7:         Let ${R^{*}}=\max\{R:S_{R}\neq\emptyset\}$.
8:         Simulate ES-ColPathRulingSet$({R^{*}})$ for all nodes $u\in S_{R^{*}}$.
9:         Let $r$ be the known ruling set node closest to $v_{\alpha}$, of color $\chi(r)$.
@@
In SearchingWalk (Algorithm 6), agents to a nearby colored node of the computed ruling set. Once at this node, the agents alternate calls to ZWalk and waiting periods, according to the color of the ruling set node. The fact that close enough ruling set nodes have distinct colors guarantees that the two agents, if they have computed a ruling set with the same distance $R\geq D$, will have different behaviors at some point. I.e., one of the agent will be performing a ZWalk while the other is not moving. The agents achieve rendezvous as this happens.
Algorithm 6 $\textnormal{{{SearchingWalk}}}(R,L,r,c)$, in which agent $\alpha$ searches in the distance $20R$-neighborhood of node $r$ close to its initial position $v_{\alpha}$
@@
3:ZWalk$(8R)$. $\triangleright$ $32R$ rounds
4:for $i$ from 00 to $4$ do $\triangleright$ iterating over the bits of $c=\chi(r)\in\{0,1\}^{5}$
@@
6:         ZWalk$(8R)$ twice. $\triangleright$ $2\times 32R$ rounds
7:     $\triangleright$ $c_{i}=0$
@@
Missing from this rough overview is notably how the argument is robust to the fact that agents can have a delay of $\tau$ between their wake-up times. In the next sections, we analyze our algorithm and prove Theorem 1.1, that is, that the agents achieve rendezvous in $O(D\log^{*}\ell_{\min})$ rounds when executing Algorithm 5. We first show that once $L$ reaches a large enough $\Theta(D)$ for one of the agents, if the two agents are too out-of-sync, they achieve rendezvous during the discovery phase of this large enough parameter $L$ (Lemma 4.3). We then show that if the searching phases of a large enough parameter $L$ of the two agents are sufficiently in-sync, the two agents achieve rendezvous during these searching phases (Lemmas 4.5, 4.6 and 4.7). The two arguments are put together in Section 4.4 to prove our main result.
We use the following notation in what follows:
  • For each $L\in\{2^{i}\mid i\in\mathbb{N}\}$, $T_{\alpha,L}$ is the round in which agent $\alpha$ starts its discovery phase with parameter $L$ (step $4$ in Algorithm 5). $T_{\beta,L}$ is the equivalent for agent $\beta$.
@@
A first remark about our algorithm is that the discovery and searching phases of our agents have fixed runtimes.
@@ 4.2 First Argument: Out-of-Sync Rendezvous
For the first inequality, first note that by definition $x(T_{\alpha,L})=x(0)=v_{\alpha}$ since the agent always returns at $v_{\alpha}$ before starting a new loop iteration. This immediately implies that $\lvert x(t)-x(0)\rvert\leq L/2$ for $t\in[T_{\alpha,L},T_{\alpha,L}+L/2]$, as agents can only cross one edge per round. For smaller $t$ ($t\in[0,T_{\alpha,L}]$, let us consider the movements done in loop iterations for smaller values of $L$. As an agent executes a ZWalk with some parameter $d$, it stays at distance $\leq d$ from the node where it started its execution. As a results, all calls to ZWalk as part of previous discovery phases have kept the agent at distance $\leq L/2$ from $x(0)=v_{\alpha}$. As for the movements done in the searching phase, the agent moves at distance at most $2{R^{*}}-1+8{R^{*}}<L$ from $v_{\alpha}$. Indeed, at most $12{R^{*}}-$ 1 are done to go to a nearby ruling set node, from which ZWalk is called with parameter $8{R^{*}}$, and ${R^{*}}\leq L/16$.
The second item simply follows from the fact that a ZWalk of parameter $L$ consists of going $L$ steps away from the starting position, then taking $2L$ steps in the other direction. Therefore, either $x(T_{\alpha,L}+L)=v_{\alpha}-L$ and $x(T_{\alpha,L}+3L)=v_{\alpha}+L$, or $x(T_{\alpha,L}+L)=v_{\alpha}+L$ and $x(T_{\alpha,L}+3L)=v_{\alpha}-L$. In either case, the entire interval $[v_{\alpha}-L,v_{\alpha}+L]$ is swept. ∎
@@
Note that $L\in[2D,4D-1]$. If $\tau\geq 5L/2$, then for all $t\leq T_{\beta,L}+L/2=T_{\alpha,L}+\tau+L/2\leq T_{\alpha,L}+3L$, $\lvert y(t)-v_{\beta}\rvert\leq L/2$. Because $\lvert v_{\alpha}-v_{\beta}\rvert=D$, this also implies that $\lvert y(t)-v_{\alpha}\rvert\leq D+L/2\leq L$ for all $t\leq T_{\alpha,L}+3L$. Since $\alpha$ explores the entire distance-$L$ neighborhood of $v_{\alpha}$ between $T_{\alpha,L}+L$ and $T_{\alpha,L}+3L$, by Proposition 3.2, the two agents achieve rendezvous before round $T_{\alpha,L}+3L$. ∎

### v2

@@ 3.3 Computing Colored Ruling Sets on Paths, and Stopping Early
Our algorithm gradually computes a ruling set $S$ of the infinite line, in the sense that we provide an algorithm in which every node on the infinite line eventually decides whether it is part of the ruling set or not, and in any given round, the set of nodes that committed to their output and joined the set $S$ is a ruling set over the nodes that committed to their output. At a high level, Algorithm˜3 partitions the nodes on the infinite line according to their IDs, and computes a ruling set over each set of nodes of this partition using Algorithm˜2 (PathRulingSet) from the previous section. The computation done by the set of nodes $V_{i}=\{v\in V:\log^{*}\ell(v)=i\}$ is well defined as it is performed by a finite set of nodes, with bounded IDs. After $\Theta(R\cdot i)$ rounds, the nodes in $V_{i}$ have all finished their execution of PathRulingSet, as did previously nodes in the sets $V_{j}$, $j<i$. Once $V_{i}$ has computed a $V_{i}$-limited ruling set $S_{i}$, we use it to extend the current ruling set $S$, that only contains nodes from $V_{<i}$ at this point. This is done greedily. We first have each node $v\in S_{i}$ join $S$ if and only if $S$ does not already contain a node at distance <R<R from $v$ on the line. A second greedy extension slightly improves the covering distance. Finally, we assign a color to the new members of the ruling set.
@@
[…]
@@
4:  Compute a $V_{i}$-limited (R,R1)(R,R-1)-ruling set with PathRulingSet.
@@
7:  Let $S_{<i}$ be the value of $S$ at this stage. $\triangleright$ $S_{<i}$ is a $V_{<i}$-limited (R,R1)(R,R-1)-ruling set
@@
10:  for 44 iterations do
11:   For each $v\in V_{i}$, compute $b_{v}=\min_{u\in S}\operatorname{dist}(v,u)$. Let $U^{\geq R}_{i}=\{v\in V_{i}:b_{v}\geq R\}$.
12:   Add to $S$ all $v\in U^{\geq R}_{i}$ s.t. $\forall u\in N^{R-1}(v)\cap U^{\geq R}_{i},(b_{u}<b_{v}\vee(b_{u}=b_{v}\wedge\ell(v)>\ell(u)))$.
@@
For any round $T$, let $U_{T}$ be the set of nodes that committed to their outputs by round $T$, $S_{T}\subseteq U_{T}$ the set of nodes that joined $S$ by round $T$, and $\chi_{T}:S_{T}\to[17]$ the colors assigned to nodes in $S_{T}$. Algorithm˜3 achieves the following:
@@
    At termination, a node $v$ outputs whether it is part of the ruling set $S$ or not, and if part of $S$, it also outputs a color between $1$ and $17$. If outside $S$, $v$ knows the members of $S$ at distance $\leq R-1$ from itself together with their colors.
@@
    $S_{T}$ is a $U_{T}$-limited (R,R1)(R,R-1)-ruling set, and $\chi_{T}$ is a proper coloring of $G^{9R-1}[S_{T}]$.
@@
We first show that Algorithm˜3 (ES-ColPathRulingSet) constructs the claimed object, before bounding its complexity.
Properties of the Ruling Set.
Recall that $S_{i}$, the output of PathRulingSet executed by nodes in $V_{i}$, is a $V_{i}$-limited (R,R1)(R,R-1)-ruling set. Consider the set $S_{<i}$, the content of $S$ before adding nodes from $V_{i}$ to it. Assume $S_{<i}$ to be a $V_{<i}$-limited (R,R1)(R,R-1)-ruling set. We show that the nodes from $V_{i}$ added to it make $S$ a $V_{\leq i}$-limited (R,R1)(R,R-1)-ruling set. By induction, starting from the trivial base case of $S_{<1}$, we get the claimed properties about our ruling set.
Every node $v\in V_{i}$ is at distance at most R1R-1 from a node $u\in S_{i}\subseteq V_{i}$. Consider the state of $S$ after adding $S^{\prime}_{i}=S_{i}\setminus N^{R-1}(S)$ to it. If $u$ was added to $S$, then $v$ remains at distance at most R1R-1 from a node in $S$. If $u$ was not added to $S$, there exists a node $w\in S_{<i}$ that joined $S$ earlier, such that $\operatorname{dist}(u,w)<R$. This implies that $v$ is at distance at most $\operatorname{dist}(v,u)+\operatorname{dist}(u,w)\leq 2R-2$ from a node in $S$. At this state, $S=S_{<i}\cup S^{\prime}_{i}$ is a $V_{\leq i}$-limited (R,2R2)(R,2R-2)-ruling set. The $4$ iterations of the inner loop that add to $S$ some of the nodes in $V_{i}$ furthest away from the current set $S$ make $S$ a $V_{\leq i}$-limited (R,R1)(R,R-1)-ruling set, from similar arguments as used in the proof of Lemma˜3.4.
@@
Complexity.
Recall that PathRulingSet (Algorithm˜2) with parameter distance $R\geq 1$ has a complexity of $O(R\log^{*}n)$ rounds when run by nodes of maximum ID $n$. Nodes in $V_{i}$ thus spend $O(R\cdot i)$ rounds computing $S_{i}$ at the beginning of the (parallel) loop in Algorithm˜3.
For every $i\in\mathbb{N}^{*}$, let $T_{i}$ be the round in which nodes in $V_{i}$ terminate in Algorithm˜3. Nodes in $V_{i}$ first compute $S_{i}$, before greedily adding nodes of $S_{i}$ to $S$. Before adding nodes of $S_{i}$ to $S$, all nodes in $V_{j},j<i$ must have terminated, which requires waiting until round $T_{i-1}$ (0 for the first set of nodes). After this amount of time, inserting the nodes of $S_{i}$ into $S$ only takes R1R-1 rounds, for each node in $S_{i}$ to learn which nodes in its (R1)(R-1)-neighborhood have joined $S$. The $4$ loop iterations in which additional nodes from $V_{i}$ are added to $S$ are also completed in $O(R)$. Finally, solving the list-coloring problem is done in $O(R\cdot i)$ rounds (Lemma˜2.1), and each node learning the output of nodes within distance $R$ takes $O(R)$ rounds.
@@ 4.1 The Algorithm
In this section, we describe our deterministic rendezvous algorithm. We actually describe an algorithm working in the setting where agents achieve rendezvous when crossing the same edge in opposite directions. By Theorem˜3.1, this implies an algorithm with similar asymptotic complexity for the setting where agents only detect each other if at the same node in a given round.
@@
The two basic building blocks of our algorithm are ZWalk and ES-ColPathRulingSet, respectively Algorithm˜4 and Algorithm˜3. ZWalk is simply a procedure in which the agent executing it scans the $2L+1$ nodes of the distance-$L$ neighborhood of the node it starts the ZWalk from, in $4L$ rounds. Note that while the agents do not have a common orientation, they have a sense of direction in that they know the port from which they enter a node. This allows the agents to keep walking in the same direction for successive steps, and to change direction, as desired. Most moves taken during our algorithm are done in executions of ZWalk. ES-ColPathRulingSet is the early-stopping algorithm for computing colored ruling sets presented in the previous section.
@@
6:  if $\exists R\in\{4^{j}\mid j\in\mathbb{N}\}$, $R\leq L/16$, s.t. $S_{R}\neq\emptyset$ then
7:   Let ${R^{*}}=\max\{R\in\{4^{j}\mid j\in\mathbb{N}\}:R\leq L/16\wedge S_{R}\neq\emptyset\}$.
[…]
9:   Let $r$ be the known colored ruling set node closest to $v_{\alpha}$, breaking ties with IDs.
@@
Our algorithm (Algorithm˜5) consists of a loop with a parameter $L$ that doubles between iterations. Inside the loop are two phases, which we call the discovery phase and the searching phase. In the discovery phase, an agent explores the distance-$L$ neighborhood around their starting position, roughly doubling how much of the network they know about. The agents use this information to compute a colored ruling set of parameter $R$, where $R\leq L/16$ is the largest power of $4$ s.t. the distance-$O(R)$ neighborhood of their starting location contains a ruling-set node. Of course, as the agents have different views, they do not necessarily end up computing a ruling-set with the same parameter $R$, and we denote by $R_{\alpha,L}$ and $R_{\beta,L}$ the distance of the ruling sets computed by agents $\alpha$ and $\beta$ in their phase $L$. The agents use their respective computed ruling set to search for one another, using procedure SearchingWalk (Algorithm˜6), which constitutes the searching phase.
In SearchingWalk (Algorithm˜6), each agent goes to a nearby colored node of the computed ruling set. Once at this node, the agents alternate calls to ZWalk and waiting periods, according to the color of the ruling set node. The fact that close enough ruling set nodes have distinct colors guarantees that the two agents, if they have computed a ruling set with the same distance $R\geq D$, will have different behaviors at some point. That is, one of the agents will be performing a ZWalk while the other is not moving. The agents achieve rendezvous as this happens.
Algorithm 6 $\textnormal{{{SearchingWalk}}}(R,L,r,c)$, in which agent $\alpha$ searches in the distance-$8R$ neighborhood of node $r$ close to its initial position $v_{\alpha}$
@@
3:ZWalk(8R)(8R). $\triangleright$ $32R$ rounds
4:for $i$ from 0 to 44 do $\triangleright$ iterating over the bits of $c=\chi(r)\in\{0,1\}^{5}$
@@
6:   ZWalk(8R)(8R) twice. $\triangleright$ $2\times 32R$ rounds
[…]
@@
4.2 Analysis Overview
Missing from the previous section’s rough overview is notably how the argument is robust to the fact that agents can have a delay of $\tau$ between their wake-up times. We now explain how, in the next sections, we analyze our algorithm and prove Theorem˜1.1, that is, that the agents achieve rendezvous in $O(D\log^{*}\ell_{\min})$ rounds when executing Algorithm˜5.
Organization of the proof.
To prove Theorem˜1.1, we first bound in Section˜4.3 how out-of-sync the two agents can be without trivially achieving rendezvous. Knowing that a delay $\tau$ that exceeds some threshold $\Theta(D)$ necessarily implies that the agents achieve rendezvous in $O(D)$ rounds, the rest of the analysis can assume some degree of synchronization between the agents.
We then argue that as the agents reach a large enough value of $L$ (of order $\Theta(D\log^{*}\ell_{\min})$) in their executions of Algorithm˜5, they compute nodes from colored ruling sets parameterized by distances of order at least $\Omega(D)$. When using such ruling set nodes in their executions of SearchingWalk (Algorithm˜6), the agents can achieve rendezvous in two different ways. If the two agents have computed nodes from distinct ruling sets (i.e., the distance parameters of their ruling sets do not match), the agent with the larger distance is guaranteed to find the other agent. This follows from the agent with the smaller distance parameter being constrained to some interval on the line, which we show gets fully explored by the agent with the larger parameter. We show this in Section˜4.4.1. The analysis for when the two agents computed nodes from the same ruling set (i.e., with the same distance parameter) is in Section˜4.4.2. There, we argue that rendezvous is easily achieved if the two agents decided to go to the same node of the ruling set, and if not, they find each other from the fact that they perform SearchingWalk from nodes that are both close enough to one another, and are colored with distinct colors. The various arguments are put together in Section˜4.5.
Notation and preliminary remarks.
We use the following notation:
  • For each $L\in\{2^{i}\mid i\in\mathbb{N}\}$, $T_{\alpha,L}$ is the round in which agent $\alpha$ starts its discovery phase with parameter $L$ (line $3$ in Algorithm˜5). $T_{\beta,L}$ is the equivalent for agent $\beta$.
@@ 4.2 Analysis Overview
A useful remark about our algorithm is that the discovery and searching phases of our agents have fixed runtimes.
@@ 4.3 First Argument: Out-of-Sync Rendezvous
For the first inequality, first note that by definition $x(T_{\alpha,L})=x(0)=v_{\alpha}$ since the agent always returns at $v_{\alpha}$ before starting a new loop iteration. This immediately implies that $\lvert x(t)-x(0)\rvert\leq L/2$ for $t\in[T_{\alpha,L},T_{\alpha,L}+L/2]$, as agents can only cross one edge per round. For smaller $t$ ($t\in[0,T_{\alpha,L}]$, let us consider the movements done in loop iterations for smaller values of $L$. As an agent executes a ZWalk with some parameter $d$, it stays at distance $\leq d$ from the node where it started its execution. As a results, all calls to ZWalk as part of previous discovery phases have kept the agent at distance $\leq L/2$ from $x(0)=v_{\alpha}$. As for the movements done in the searching phase, the agent moves at distance at most $2{R^{*}}-1+8{R^{*}}<L$ from $v_{\alpha}$. Indeed, at most $2{R^{*}}-1$ are done to go to a nearby ruling set node, from which ZWalk is called with parameter $8{R^{*}}$, and ${R^{*}}\leq L/16$.
The second item simply follows from the fact that a ZWalk of parameter $L$ consists of going $L$ steps away from the starting position, then taking $2L$ steps in the other direction. Therefore, either $x(T_{\alpha,L}+L)=v_{\alpha}-L$ and $x(T_{\alpha,L}+3L)=v_{\alpha}+L$, or $x(T_{\alpha,L}+L)=v_{\alpha}+L$ and $x(T_{\alpha,L}+3L)=v_{\alpha}-L$. In either case, the entire interval $v_{\alpha}-L,v_{\alpha}+L]$ is swept. ∎
@@
Note that $L\in[2D,4D-1]$. If $\tau\geq 5L/2$, then for all $t\leq T_{\alpha,L}+3L\leq T_{\alpha,L}+\tau+L/2=T_{\beta,L}+L/2$, $\lvert y(t)-v_{\beta}\rvert\leq L/2$ by Proposition˜4.2. Because $\lvert v_{\alpha}-v_{\beta}\rvert=D$, this also implies that $\lvert y(t)-v_{\alpha}\rvert\leq D+L/2\leq L$ for all $t\leq T_{\alpha,L}+3L$. Since $\alpha$ explores the entire distance-$L$ neighborhood of $v_{\alpha}$ between $T_{\alpha,L}+L$ and $T_{\alpha,L}+3L$ by Proposition˜4.2, the two agents achieve rendezvous before round $T_{\alpha,L}+3L$ by Proposition˜3.2. ∎